# MG-SLAM: Structure Gaussian SLAM with Manhattan World Hypothesis

### v2

@@ 3.3.1 Scene representation
where $\mu_{\mathcal{G}}$ and $\Sigma_{\mathcal{G}}$ are mean and covariance of the Gaussian. $J$ is the Jacobian of the linear approximation of $\pi(\cdot)$. For each pixel ${\rm pix}\in\mathcal{P}$, the influence of $N$ Gaussians on this pixel can be combined by sorting the Gaussians in depth order and performing front-to-back volume rendering using the Max volume rendering formula [29]:
$\mathcal{P}_{\rm pix}=\sum_{i\in N}p_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),~{}\mathcal{D}_{\rm pix}=\sum_{i\in N}z_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (9)
where $\mathcal{P}_{\rm pix}$ and $\mathcal{D}_{\rm pix}$ represent the pixel-wise color ($\mathcal{C}_{\rm pix}$ and $\mathcal{S}_{\rm pix}$) and depth respectively. $z$ is the distance to the mean of the Gaussian $\mathcal{G}$ along the camera ray.
3.3.2 Mapping loss
The scene representation is optimized based on the selected keyframes obtained from the tracking system. Specifically, we minimize the following photometric residual:
$\mathcal{L}_{\rm pix}=\left\|\mathcal{C}_{\rm pix}-\mathcal{C}^{\rm GT}_{\rm pix}\right\|+\lambda_{\mathcal{S}}\left\|\mathcal{S}_{\rm pix}-\mathcal{S}^{\rm GT}_{\rm pix}\right\|+\lambda_{\mathcal{D}}\left\|\mathcal{D}_{\rm pix}-\mathcal{D}^{\rm GT}_{\rm pix}\right\|$ (10)
In this context, Gaussians are optimized to adjust their optical and geometrical parameters to closely replicate the observed scene with fine-grained details. Furthermore, to impose constraints on the scene utilizing the identified line features that precisely delineate the lines and potential edges in the current view, an additional line residual is introduced for pixels associated with line features to enhance the map’s accuracy. Consequently, the aggregate loss for mapping is determined as follows:
@@ 3.4 Structure optimization
In our optimization process, we focus particularly on refining the representation of expansive surfaces that define the space, such as the floor and ceiling. By applying the Manhattan World assumption, we introduce the planar hypothesis surfaces that are informed by the structural regularity of the extracted 3D line features. These constraints are employed to identify surfaces that are not adequately represented and to interpolate new Gaussians for densifying or filling in gaps with textures that are reasonably predicted. Specifically, our method consists of the following steps:
@@
Calibration  Since the SLAM system relies on the initial camera pose as a reference frame, the reconstructed scenes usually do not satisfy our orthogonality assumption. As a result, we calibrate the reconstructed scene by applying a calibrating matrix $K$ to the coordinates and rotations of the Gaussians. This matrix $K$ is derived from clustering the directions of 3D line segments that are presumed to align with the scene axes.
Surface interpolation  After aligning the structured surface boundaries to orthogonal directions, they extend across the $xy\rightarrow\mathbb{R}^{2}$ plane. We use the calibrated line segments to outline the rectangular boundary of the hypothesized plane, capitalizing on the dense line features commonly found at the corners of scenes. The chosen target Gaussians representing the surface are then projected from their 3D form onto the 2D hypothesis $xy$ plane. Then we apply a density threshold to a predefined grid using the density function (Equation 13) to detect any holes or gaps on the surface. New Gaussians are generated at the identified gaps that fall below a density threshold. We employ a PointNet++ model ([34]), fitting it to the presented surface to learn the color and texture patterns. This model then interpolates the texture color of the new Gaussians based on their spatial properties.
@@
Through this method, we aim to seamlessly combine the detailed representation provided by Gaussians with the need for explicit scene geometry that accounts for unseen views, thereby achieving a more comprehensive and accurate reconstruction of indoor environments. The detailed explanation of the algorithms is outlined in Section B.2.
4 Experiment
4.1 Experimental setup
@@
Datasets  We evaluate our method using two datasets: Replica [38], a synthetic dataset, and ScanNet [6], a challenging real-world dataset. For Replica [38], the ground-truth camera pose and semantic maps are obtained through Habitat simulation [37]. In the case of ScanNet [6], the ground-truth camera poses are derived using BundleFusion [7].
@@
Metrics  To assess the quality of the reconstruction, we employ metrics such as PSNR, Depth-L1, SSIM, and LPIPS. For evaluating camera pose, we use the average absolute trajectory error (ATE RMSE). The real-time processing capability, essential for SLAM systems, is measured in frames per second (FPS).
@@
Baselines  We evaluate our tracking and mapping results against state-of-the-art methods, including NeRF-based approaches such as NICE-SLAM [51], ESLAM [18], and Point-SLAM [36], as well as recent Gaussian-based methods like SplaTAM [19] and MonoGS [28]. The results for MonoGS [28] are obtained using its RGB-D mode for fair comparison.
Ablation  The ablation study demonstrating the effectiveness of each proposed component is detailed in Appendix C.
4.2 Evaluation of localization and reconstruction
@@
Figure 4: Qualitative comparison of our method and the baselines for novel-view synthesis on the ScanNet dataset [6]. The outcomes show that our method provides more robust and fine-grained reconstructions in real-world complex scenes.
In Figure 4, we visualize the novel-view synthesis results for MG-SLAM, comparing it against baseline methods in real-world scenes. Our method exhibits robust and high-fidelity reconstruction capabilities. In comparison to Point-SLAM [36], MG-SLAM offers complete scene reconstructions and finer texture details, benefiting from the use of Gaussians to handle complex geometries. Among Gaussian-based methods, SplaTAM [19] struggles in large multi-room scenes with complex camera loops due to its lack of bundle adjustment in the tracking system. Similarly, MonoGS [28] generally delivers reliable quality but struggles with object-level reconstruction drift. In contrast, MG-SLAM excels with robust tracking and effective bundle adjustment incorporating line segments, enabling superior detailed reconstructions even surpassing the ground-truth mesh derived from offline methods.
We provide quantitative assessments of reconstruction quality using the ScanNet dataset [6] in Table 1 and the Replica-V1 dataset [38] in Appendix D.1. Our approach delivers SOTA results, outperforming other Gaussian-based methods by a notable 4dB in PSNR on both datasets. The tracking evaluation results are shown in Table 2. Our method remarkably reduces the ATE RMSE (cm) error, achieving 50% improvements over Gaussian baselines. Moreover, benefiting from our line fusion strategy, MG-SLAM also shows superior tracking performance over traditional SLAM systems [31, 2, 14], as detailed in Section D.3.
@@
Table 1: Quantitative comparison of our method and the baselines in training view rendering on the ScanNet dataset [6]. Our method demonstrates SOTA performances among three metrics.
@@
Table 2: Quantitative comparison of our method and the baselines in terms of ATE [cm] on the ScanNet dataset [6]. Our method demonstrates SOTA performances in most of presented scenes.
  Methods Avg. $0$ 0059 0106 0169 0207
ESLAM [18] 7.68 8.47 8.70 7.58 7.45 6.20
Point-SLAM [36] 11.68 10.24 7.81 8.65 22.16 9.54
SplaTAM [19] 12.04 12.83 10.10 17.72 12.08 7.46
MonoGS [28] 13.86 15.94 13.03 19.44 10.44 10.46
Ours 6.77 5.95 6.41 8.07 7.29 6.14
4.3 Evaluation on large-scale dataset
Figure 5: Qualitative comparison of our method and the Gaussian SLAM baselines for novel-view synthesis on the Replica Apartment dataset [38]. More detailed figures are available in Section D.2.
Table 3: Quantitative comparison of our method and the Gaussian baselines in training view rendering on the Replica Apartment dataset [38]. Our method demonstrates SOTA performances.
  Methods Metrics Avg. apart.0 apart.1 apart.2 frl_apart.0 frl_apart.4

### v3

@@ III-C1 Scene Representation
where $\mu_{\mathcal{G}}$ and $\Sigma_{\mathcal{G}}$ are mean and covariance of the Gaussian primitives. $J$ is the Jacobian of the linear approximation of $\pi(\cdot)$. For each pixel ${\rm pix}\in\mathcal{P}$, the influence of $N$ Gaussians on this pixel can be combined by sorting the Gaussians in depth order and performing front-to-back alpha-blending:
$\mathcal{P}_{\rm pix}=\sum_{i\in N}f_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),~{}\mathcal{D}_{\rm pix}=\sum_{i\in N}z_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (10)
where $\mathcal{P}_{\rm pix}$ and $\mathcal{D}_{\rm pix}$ represent the pixel-wise appearance ($\mathcal{C}_{\rm pix}$ for color and $\mathcal{S}_{\rm pix}$ for semantic features) and depth respectively. $z$ is the distance to the mean of the Gaussian $\mathcal{G}$ along the camera ray.
III-C2 Mapping Loss
The scene representation is optimized using keyframes obtained from the tracking system by minimizing the photometric residual:
$\mathcal{L}_{\rm pix}=\left\|\mathcal{C}_{\rm pix}-\mathcal{C}^{\rm GT}_{\rm pix}\right\|$ $+\lambda_{\mathcal{S}}\left\|\mathcal{S}_{\rm pix}-\mathcal{S}^{\rm GT}_{\rm pix}\right\|$
$+\lambda_{\mathcal{D}}\left\|\mathcal{D}_{\rm pix}-\mathcal{D}^{\rm GT}_{\rm pix}\right\|$ (11)
In this context, Gaussian primitives are optimized to adjust their optical and geometrical parameters to closely replicate the observed scene with fine-grained details. Furthermore, to impose constraints on the scene utilizing the identified line features that precisely delineate the lines and potential edges in the current view, an additional line residual is introduced for pixels associated with line features to enhance the map’s accuracy. Consequently, the aggregate loss for mapping is determined as follows:
@@ III-D Structure Optimization
In our optimization process, we focus particularly on refining the representation of expansive surfaces that formulate the indoor scene, such as the floor and ceiling. By applying the MW assumption, we introduce the planar hypothesis surfaces that are informed by the structural regularity of the extracted 3D line features. These constraints are employed to identify surfaces that are not adequately represented and to interpolate new Gaussians for densifying or filling in gaps with textures that are reasonably predicted.
III-D1 Gaussian Density
In our Gaussian map, we define the density of the Gaussian $\nu:\mathbb{R}^{3}\rightarrow\mathbb{R}_{+}$ as the sum of the Gaussian values weighted by their alpha-blending coefficients at any given grid points p as:
where the $\mu_{\mathcal{G}_{i}},\Sigma_{\mathcal{G}_{i}},\alpha_{\mathcal{G}_{i}}$ are the means, covariances, and alpha-blending coefficients of the Gaussians $\mathcal{G}$. To simplify the calculation, $\nu(p)$ can be approximated following [31] as:
where $\mathcal{G}^{*}$ is the nearest Gaussian that contributes most to $p$. This Gaussian density function $\nu^{*}(\cdot)$ facilitates the subsequent identification of under-represented regions on the planar surfaces and the mesh generation process.
III-D2 Map Calibration
@@
Since the SLAM system relies on the initial camera pose as a reference frame, the reconstructed scenes usually do not satisfy the orthogonality assumption of the Manhattan World. To overcome this issue, we calibrate the reconstructed scene by applying a calibrating matrix $K$ to the coordinates and covariance of the Gaussians. This matrix $K$ is derived from clustering the directions of 3D line segments that are presumed to align with the scene axes.
After aligning the structured surface boundaries to orthogonal directions, they extend across the $xy\rightarrow\mathbb{R}^{2}$ plane. We use the calibrated line segments to outline the rectangular boundary of the hypothesized planes, capitalizing on the dense line features commonly found at the corners of scenes. Algorithm 1 presents the detailed implementation.
1:Line segments $L\in\mathbb{R}^{3}$ and the reconstructed Gaussian map $\mathcal{G}$
3:for each line segment $\ell$ in $L_{\rm filtered}$ do
4:    $\{direction\}\leftarrow(\ell.end-\ell.start).normalized()$
5$\triangleright$ Add direction vector
6:end for
7:$\vec{x},\vec{y},\vec{z}\leftarrow k\_means(\{direction\},k=3)$triangleright Find axes directions
8:$K\leftarrow\mathbb{I}\in\mathbb{R}^{4}$triangleright Define calibration matrix
[…]
12:for each $\mathcal{G}_{i}$ in $\mathcal{G}$ do
13:    $\mu_{\mathcal{G}_{i}}\leftarrow K\cdot\mu_{\mathcal{G}_{i}}$triangleright Calibrate coordinates of Gaussian
14:    $\gamma_{\mathcal{G}_{i}}\leftarrow K\cdot\gamma_{\mathcal{G}_{i}}$triangleright Calibrate rotation of Gaussian
15:end for
16:for each line segment $\ell$ in $L$ do
17:    $\ell\leftarrow K\cdot\ell$triangleright Calibrate line segment
18:end for
$\mathcal{B}\leftarrow$ select maximum and minimum $x$ and $y$ for endpoints in $L$
20$\triangleright$ Define surface boundary
21:Output: boundary of surface $\mathcal{B}\in\mathbb{R}^{2}$
Algorithm 2 The pseudo-code for scene completion
3:for each structure surface $\mathcal{G}_{i}$ in $\mathcal{G}(\mathcal{M}_{tg})$ do
4:    $\mathcal{G}_{i}\leftarrow DBSCAN.fit(\mathcal{G}_{i})$
5$\triangleright$ Apply DBSCAN [71] to eliminate outliers
6:    $\bar{z}\leftarrow avg(z)$ for $z\in\mu_{\mathcal{G}_{i}}$
7:    $\Pi\leftarrow grid(\mathcal{B})$triangleright Create the 2D grid of hypothesis surface
8:    $\mathcal{D}_{i}\leftarrow$ compute the Gaussian density for each point of the grid in $\Pi$.
9:    if $\mathcal{D}_{i}\leq\mathcal{T}_{density}$ then
10:        $\mu_{\rm pred}\leftarrow$ generates new Gaussians centers
11:    end if
13:    for each batch $B_{i}$ in $B^{*}_{\mathcal{G}_{i}}$ do
15:    end for
16:    $\mathcal{C}_{\rm pred}\leftarrow\mathcal{F}(\mu_{pred})$triangleright Predict the color of new Gaussians
17:    $\mathcal{G}_{\rm pred}\leftarrow$ formulate new Gaussians using $\{\mu_{\rm pred}\}$ and $\{\mathcal{C}_{\rm pred}\}$
18:end for
19:$\mathcal{G}\leftarrow\mathcal{G}_{\rm pred}$triangleright Update the overall scene
20:Output: structure optimized $\mathcal{G}$
III-D3 Surface Interpolation
Algorithm 2 shows our surface interpolation strategy. Specifically, we utilize the semantic information, incorporated in Equation 11, for identifying the planar surfaces such as floors and ceilings. The chosen target Gaussians are then projected from the 3D space onto the 2D hypothesis $xy$ plane. Subsequently, we apply a density threshold to the sampling density function, defined in Equation 14, to detect potential holes or gaps on the surface. New Gaussian primitives are generated at the identified gaps that fall below a density threshold. Finally, we train a PointNet++ model [72], fitting it to the presented surface to learn the presented color patterns and interpolate the texture color of the new Gaussians based on their spatial correlation. We empirically found that the PointNet++ network, which incorporates spatial correlation, is sufficiently effective in predicting the color patterns of new primitives on the gaps of structured surfaces. Given that these planar surfaces typically exhibit less texture than specific objects, employing more sophisticated models does not provide substantial benefits but increases computational expenses.
@@ III-D3 Surface Interpolation
Through this method, we aim to seamlessly combine the detailed representation provided by Gaussians with the need for explicit scene geometry that accounts for unseen views, thereby achieving a more comprehensive and accurate reconstruction of planar surfaces in indoor environments.
III-E Mesh Generation
We employed the surface extraction method proposed by [31] with a novel normal regularization term during the scene refinement. To compress the Gaussians to align closely with the surface, [31] forces the Gaussians to be flat by setting one of their scaling factors nearly to zero. Consequently, $\nu^{*}(p)$ can be further simplified to $\hat{\nu}^{*}(p)=\alpha_{\mathcal{G}^{*}}\mathrm{exp}\left(-{2s^{-2}_{\mathcal{G}^{*}}}\langle p-\mu_{\mathcal{G}^{*}},n_{\mathcal{G}^{*}}\rangle^{2}\right)$ where $s_{\mathcal{G}}$ represents the smallest scaling factor and $n_{\mathcal{G}}$ its corresponding normal. Here, $\alpha_{\mathcal{G}^{*}}=1$ is set to avoid semi-transparent Gaussians.
Moreover, to derive smooth surfaces, particularly in flat and textureless floor areas, we adopt the MW hypothesis, which assumes that planar floors are orthogonal to the vertical axis. We enforce this constraint by introducing a normal regularization term for the estimated SDF:
In this context, the projected points $x_{\rm 2D}$, which pass through the camera ray and fall within the floor area, are encouraged to align their normal to the ideal vector $\hat{n}=\langle 0,0,1\rangle$. The set $\mathcal{P}_{\text{floor}}$ consists of image pixels identified as the floor using semantic segmentation.
IV Experiment
IV-A Experiment Settings
Implementation Details Our tracking system is implemented in C++, while the mapping system uses Python3 and CUDA C. The experiments were conducted using a single RTX A100-80GB GPU and 24-core AMD EPYC 7402 processor. We use Adam optimizer for Gaussian representation optimization and network training. The hyperparameters for 3D Gaussian splitting [26] are the same as the original paper. Table I presents the values for hyperparameters in our system. For Gaussian insertion, in each keyframe, we sampled $\mu_{\mathcal{G}}$ in Equation 9 from the rendered depth map $\mathcal{D}_{\rm pix}$ following the distribution of $\mathcal{N}(\mathcal{D}_{\rm pix},0.2\sigma_{\mathcal{D}})$. For uninitialized areas, we initialize Gaussians by sampling from $\mathcal{N}(\bar{\mathcal{D}}_{\rm pix},0.5\sigma_{\mathcal{D}})$ where $\bar{\mathcal{D}}_{\rm pix}$ is the mean value. For pruning, we remove the Gaussians with opacity less than 0.6. Our system incorporates a segmentation loss introduced in [23], and we leverage this semantic information to effectively extract structured surfaces, such as floors and ceilings, from the reconstructed scenes. We subsequently train a PointNet++ [72] network for each structure surface and interpolate the missing geometry by predicting the color of inserted Gaussians.
@@ IV-A Experiment Settings
Datasets We evaluate our method using two datasets: Replica [73], a synthetic dataset, and ScanNet [27], a challenging real-world dataset. The large-scale Replica Apartment dataset used in our experiments was released by Tandem [74]. For the Replica dataset [73], the ground-truth camera pose and semantic maps are obtained through Habitat simulation [75]. In the case of the ScanNet dataset [27], the ground-truth camera poses are derived using BundleFusion [76]. Moreover, we further validate our method on a private data trajectory collected using our physical platform.
@@
Metrics To assess the quality of the reconstruction, we employ metrics such as PSNR, SSIM, and LPIPS. For evaluating camera pose, we use the average absolute trajectory error (ATE RMSE). The real-time processing capability, essential for SLAM systems, is measured in frames per second (FPS). Best results are shaded as first, second, and third.
@@
Baselines We evaluate our tracking and mapping results against state-of-the-art methods, including NeRF-based approaches such as NICE-SLAM [1], Co-SLAM [3], ESLAM [4], and Point-SLAM [6], as well as recent Gaussian-based methods including SplaTAM [20], MonoGS [21], Photo-SLAM [24], and RTG-SLAM [39]. The results for MonoGS [21] were obtained using its RGB-D mode.
TABLE I: MG-SLAM hyperparameters
Symbol Explanation Value
$N_{p}$ number of point features in tracking 4e3
$N_{\ell}$ number of line segments in tracking 200
$r_{d}$ downsample ratio for Gaussian initialization 8
$n_{map}$ mapping iteration 150
$SH_{degree}$ degree of the spherical harmonics 0
$l_{feat}$ learning rate for SH features 2.5e-3
$l_{opacity}$ learning rate for opacity 0.05
$l_{scale}$ learning rate for scaling 1e-3
$l_{rotat}$ learning rate for rotation 1e-3
$\lambda_{\mathcal{D}}$ weighting term for depth loss in mapping 0.10
$\lambda_{\mathcal{S}}$ weighting term for semantic loss in mapping 0.10
$\lambda_{\ell}$ re-weighting term for line feature loss 0.25
$\mathcal{T}_{\ell}$ line segment filtering threshold 0.08
$\mathcal{T}_{density}$ density threshold for hole detection 0.90
$b_{\mathcal{G}}$ batch size for PointNet++ [72] training 1e4
$n_{e}$ number of epoch to train PointNet++ [72] 25
IV-B Evaluation on Replica-V1 Dataset
TABLE II: Quantitative comparison of our method and the baselines in training view rendering on the Replica-V1 dataset [73]. Best results are shaded as first, second, and third.
[…]
TABLE III: System comparison in terms of tracking, mapping, rendering FPS, and memory usage between our method and the neural dense baselines on the Replica dataset [73]. The values represent the average outcomes across 8 scenes. Note that RTG-SLAM [39] and Photo-SLAM [24] incorporate ORB-SLAM2 [45] and ORB-SLAM3 [77] for tracking.
  Methods Track. ATE [cm]$\downarrow$ Track. FPS [f/s]$\uparrow$ Map. FPS [f/s]$\uparrow$ Render. FPS [f/s]$\uparrow$ SLAM FPS [f/s]$\uparrow$ Param. Size [mb]$\downarrow$
NeRF Co-SLAM [3] 1.12 10.2 10.0 35 9.26 8.2
ESLAM [4] 0.63 9.92 2.23 2.5 1.82 73.7
Point-SLAM [6] 0.54 0.42 0.06 1.3 0.05 69.6
SplaTAM [20] 0.55 0.35 0.22 122 0.14 277.1
3DGS MonoGS [21] 0.58 2.57 3.85 642 1.80 42.2
Photo-SLAM [24] 0.59 38.72 30.04 765 16.46 59
RTG-SLAM [39] 0.49 45.36 18.67 692 12.65 71
Ours 0.45 35.32 3.96 324 3.76 78.5
In line with conventional evaluations from previous studies, we performed quantitative comparisons of training view rendering on the Replica-V1 dataset [73], which comprises 8 single-room scenes. As presented in Table II, our method demonstrates superior rendering quality. This improvement stems from the use of a denser primitive distribution, as adopted in MonoGS [21], leading to higher PSNR values, and the integration of the segment reconstruction loss, which contributes to a competitive structural similarity score.
TABLE IV: Quantitative comparison of our method and the baseline approaches in training view rendering on the Replica Apartment dataset [73]. The underline indicates that relocalization was triggered due to accumulated tracking errors. The dash indicates system failures.
  Methods Metrics Avg. ap.0 ap.1 ap.2 frl.0 frl.4
[…]
Additionally, Table III provides comparisons of tracking accuracy and system efficiency. By leveraging line features, our approach achieves state-of-the-art tracking performance compared to baseline methods, further contributing to optimal map reconstruction quality. Regarding system efficiency, we assessed the frame rate of each component and memory usage. Unlike systems employing frame-to-frame tracking [3, 4, 6, 20, 21], which exhibit significantly low tracking speeds, methods such as Photo-SLAM [24], RTG-SLAM [39], and ours achieve notable advantages in tracking performance with exceptional frame rates. This is attributed to the incorporation of traditional feature-based optimization, which avoids iterative rendering and photometric loss computations for camera pose optimization—a common bottleneck in neural dense systems—thereby remarkably enhancing tracking speed. For mapping, compared to Photo-SLAM [24] and RTG-SLAM [39], which utilize relatively sparse primitive representations, our method opts for a denser map to improve reconstruction quality. While this results in slower mapping speeds and moderate memory usage, it offers a balanced trade-off for higher reconstruction fidelity.
Figure 3: The novel view sythesis of the scene apartment_0 from the Replica Apartment dataset [73]. The top left shows the line segments extracted in 3D space. The bottom left illustrates the overall reconstructed scene.
IV-C Evaluation on Replica-Apartment Dataset
TABLE V: Quantitative comparison of our method with traditional, NeRF-based, and Gaussian-based RGB-D SLAM systems in terms of ATE [cm] on the Replica Apartment dataset [73]. The underlined values indicate that relocalization was necessary due to loss of tracking.
Methods Avg. ap.0 ap.1 ap.2 frl.0 frl.4
Trad. ORB-SLAM2 [45] 5.06 6.85 9.83 6.71 1.29 1.00
ORB-SLAM3 [77] 4.41 5.39 7.98 6.31 1.09 1.27
PLVS [13] 3.72 7.06 4.67 4.04 1.45 1.39
NeRF Co-SLAM [3] 4.14 5.30 4.78 6.31 2.01 2.30
ESLAM [4] 5.54 9.22 7.09 7.01 3.33 1.07
Point-SLAM [6] 5.88 6.39 17.28 3.54 1.02 1.19
SplaTAM [20] 6.97 25.14 15.67 6.17 2.74 3.31
3DGS MonoGS [77] 6.55 20.69 9.51 7.57 2.98 6.14
Photo-SLAM [24] 4.41 5.39 7.98 6.31 1.09 1.27
RTG-SLAM [39] 5.06 6.99 9.09 6.90 1.33 0.98
Ours 2.77 5.03 3.87 3.15 0.92 0.88
To evaluate the robustness of the system in large-scale indoor environments, we evaluate MG-SLAM on the Replica Apartment dataset [73]. This dataset contains extensive multi-room scenes, complex object geometries, and looping trajectories across rooms. Table IV presents the rendering quality of our method compared to both NeRF-based and Gaussian-based approaches over five selected scenes. MG-SLAM shows notable improvements, particularly achieving a 7dB improvement in the apartment_0 scene over Gaussian SLAM systems. This optimal performance is largely attributed to the inclusion of the fused line segments, which lays a solid foundation for loop closure and pose optimization. Compared to the NeRF-based Point-SLAM [6], which utilizes additional neural point clouds for robust tracking and fine-grained mapping, our method delivers superior averaged reconstruction outcomes while maintaining optimal real-time processing capabilities, achieving a system frame rate 50 times faster.
Table V presents the quantitative evaluations of tracking ATE [cm] on the apartment scenes. Our method achieves state-of-the-art tracking accuracy compared to both frame-to-frame and feature-based tracking systems. Specifically, ORB-based systems [45, 77, 24, 39] struggle in textureless planar environments, such as the stairs or corridors in apartmen_0 and apartment_1, leading to tracking failures, frequent relocalizations, and errors in the reconstructed scenes. Leveraging the segment fusion strategy, which provides robust line features, our method also outperforms PLVS [13] under scenarios involving extensive camera movement and rotation which are commonly encountered in the apartment dataset.
Moreover, we conducted qualitative comparisons on novel-view synthesis against Gaussian-based approaches [20, 21, 24] as shown in Figure 3. These comparisons highlight the superior performance of our method over baseline approaches in large-scale complex indoor environments.
IV-D Evaluation on ScanNet Dataset
@@
Figure 4: Qualitative comparison of our method and the baselines for novel-view synthesis on the ScanNet dataset [27]. The outcomes show that our method provides more robust and fine-grained reconstructions in real-world complex scenes compared to current NeRF-based and Gaussian-based approaches.
@@
TABLE VI: Quantitative comparison of our method and the baselines in training view rendering on the ScanNet dataset [27].